To solve quation -21=7(p-10) which operation should be applied first.
I need help on my hw please thankyou.

Answer by ewatrrr(24785)

 
Hi,
To solve quation -21=7(p-10) which operation should be applied first.
Note: Rule of thumb is to first apply the Distributive Property.
-21 = 7(p-10) |Distributing the 7
-21 = 7p - 70
49 = 7p
7 = p